### v1

@@ 5 Conclusion
In this review, we explored the potential safety issues of AI in healthcare, highlighting AI’s rapid advancements alongside the significant safety challenges that impede its broader adoption. We identified key concerns related to AI reliability, including data harmonization, model calibration, generalization, and biases, as well as alignment issues, such as ensuring AI adheres to human-defined objectives. Furthermore, we delved into the specific risks posed by large language models (LLMs) in medical applications, such as hallucinations, privacy concerns, and difficulties in processing complex logic. Ultimately, while AI holds transformative potential for healthcare, addressing these safety risks is crucial for ensuring the ethical, reliable, and secure deployment of AI technologies in real-world medical settings.
6 Acknowledgements

### v2

@@ CONCLUSION
In this review, we discussed the potential safety issues of AI in healthcare, highlighting LLMs’ rapid advancements alongside the significant safety challenges that impede its broader adoption. We identified key concerns related to AI reliability, including data harmonization, model calibration, generalization, biases, and adaption in new population, as well as alignment issues, such as ensuring AI adheres to human-defined objectives. Furthermore, we delved into the specific risks posed by LLMs in medical applications, such as hallucinations, privacy concerns, and difficulties in processing complex logic. Ultimately, while AI holds transformative potential for healthcare, addressing these safety risks and establish trust among the general public is crucial for ensuring the ethical, reliable, and secure deployment of AI technologies in real-world medical settings.
RESOURCE AVAILABILITY
Lead contact
Requests for further information and resources should be directed to and will be fulfilled by the lead contact, Dianbo Liu (dianbo@nus.edu.sg).
Materials availability
This study did not generate new materials.
  • All the data and papers reported in this review will be shared by the lead contact upon request.
  • This paper does not report original code.
  • Any additional information required to reanalyze the papers reported in this paper is available from the lead contact upon request.
ACKNOWLEDGMENTS
@@
DECLARATION OF GENERATIVE AI AND AI-ASSISTED TECHNOLOGIES
All the scientific contents in this work including but are not limited to listed facts, explanation, views, ideas and citations are written by authors of this article. During the preparation of this work the ChatGPT-4o was used in order to polish and refine languages especially the grammars. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.